What is the theoretical TOTAL yield of ATP per glucose molecule in the entire Cellular Respiration process?
Answer #1

Answer:

36 ATP

Explanation:

Glycolysis: (6ATP)

Net ATP gain = 2ATP

NADH = 2*2 = 4ATP

Pyruvate oxidation: (6ATP)

Two NADH = 2*3 = 6ATP

Citric acid cycle: (24ATP)

Substrate level phosphorylation= 2ATP

6 NADH = 6*3 = 18ATP

2 FADH2 = 2 * 2 = 4 ATP

Total ATP produced in aerobic cellular respiration = 6 (glycolysis)+6(pyruvate oxidation)+24(Citric acid cycle) =36 ATP